OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? Answer 3 0 6 -2

OpenStudy (anonymous):

Answer: y = 0 \[-x + y + 3x - 8y = - 3 + 9\] \[2x = 7y + 6\] \[x = (7y + 6) \div 2\] \[(7y + 6) \div 2 - y = 3\] \[5y = 0\] \[y = 0\] Hope this has helped you :)

OpenStudy (anonymous):

Thank you! do you think you can help me with this one also? Solve the system 3x - 5y = 8 and 4x - 7y = 10. (1, -1) (2, 6) (6, 2) no solution infinitely many solutions

OpenStudy (anonymous):

Your welcome. 4(3x-5y=8) 12x-20y=32 3(4x-7y=10) 12x-21y=30 subtract y=2 Sub for y and you get x=6 So it would be (2,6) :)

OpenStudy (anonymous):

your a life saver! lol!

OpenStudy (anonymous):

Ha your welcome! :)